What is the period T of a planet which radius is as twice as of Earth when it completes one revolution in 875 days? ( Earth radius = 6.38 x 105 )


1
Expert's answer
2021-03-08T08:11:19-0500

Answer

According to kepler law

Relationship between time period and radius is given

"T^2\propto R^3"

Therefore time period of planet is given

"T_p=T_e\times(\frac{R_p}{R_e}) ^{\frac{3}{2}}"

"T_p=875\times(\frac{2R_e}{R_e}) ^{\frac{3}{2}}\\=2474.5day"